OpenStudy (anonymous):

solve for x 5x^2=-28x-15

OpenStudy (anonymous):

5x^2-28x-15=0 x=7/10 or x=49/10

hero (hero):

5x^2 = -28x - 15 5x^2 + 28x + 15 = 0 5x^2 + 25x + 3x + 15 = 0 5x(x+5)+3(x+5) = 0 (x+5)(5x+3) = 0 x = 5, -3/5